OpenStudy (anonymous):

-x+2y=8 find the value for x when y=-6

OpenStudy (anonymous):

-x+2y=8 -x+3(-6)=8 -x-18=8 18-8=x 10=x So, x=10. You just plug in -6 in for Y and you get 10 for X. Just solve as a normal equation.